Sasha owns two investments, A and B, that have a combined total value of 52,500 dollars. Investment A is expected to pay 23,900 dollars in 7 year(s) from today and has an expected return of 4.77 percent per year. Investment B is expected to pay X in 4 years from today and has an expected return of 14.86 percent per year. What is X, the cash flow expected from investment B in 4 years from today?
Answer #1

Combined Present Value of Investment A and Investment B = $52,500

Investment A is expected to pay 23,900 dollars in 7 year(s) from today

Interest Rate = 4.77%

Calculating Present Value of Investment A,

Present Value = FV/(1+r)t

Present Value = $17,248.02

Present Value of Investment B = PV of Portfolio - PV of Investment A

PV of Investment B = 52500 - 17248.02

PV of Investment B = $35,251.98

Period = 4 years

Interest Rate = 14.86%

FV = PV*(1+r)t

FV = $61,356.1

So investment B will pay $61,356.1 in Year 4